The work done by the electric force to move a charge q= 3x10-7 C is 900 J. what is the potential difference that the charge moves through?


1
Expert's answer
2020-08-17T09:27:48-0400

The work done by the electric force to move a charge

"W=qV"

The potential difference

"V=\frac{W}{q}=\frac{900\:\rm J}{3\times 10^{-7}\:\rm C}=3\times 10^9\:\rm V=3\: GV"